OpenStudy (anonymous):

how do I solve x + y = 0 and x - y + 2 = 0

OpenStudy (kirbykirby):

Solve for x in the first equation (i.e isolate it): \(x+y=0 \\ x = 0 - y \\ x = -y \) Since x = -y in this equation, plug in x = -y into the second equation: \(x - y + 2 =0 \\ (-y) - y + 2 = 0 \\ -2y + 2 = 0 \\ -2y = 0 - 2 \\ -2y = -2 \\ y = 1 \) Now you know the value of y, so substitute y = 1 into any equation to find the value of x: I will do this with the first equation: \(x + y = 0\\ x + 1 = 0\\ x = -1\) So: \(\boxed{x=-1, y=1}\)